### v1

# Contextuality of the probability current in quantum mechanics

Abstract
We revisit an argument proposed by Hardy [1] concerning local realistic theories, but in terms of probability currents in standard quantum mechanics and of the trajectories obtained from its flux lines. We emphasize surprising properties of the trajectories in configuration space, in particular their (quasi) discontinuous variations when the context (the experimental setup) is changed. This occurs in both Galilean and Einsteinian relativity. In the latter case, discontinuous variations also appear without changing the setup, when the Lorentzian reference frame is changed.
The purpose of this article is to discuss the properties of the probability current in relativistic quantum mechanics. Within Galilean relativity, one can use the Schrödinger equation, which implies the existence of a local conservation law of the probability containing a current $\bm{J}(\bm{r}_{1},\bm{r}_{2},\ldots,\bm{r}_{N},)$; this current is defined in the configuration space of the physical system (the space associated with the $N$ spatial coordinates of its particles). One can then build trajectories, which are at each point tangent to $\bm{J}$. When a different Galilean reference frame is used, the components of $\bm{J}$ are transformed as those of the current of a classical fluid. It is then possible to develop pictorial views of the motion of the probability fluid and, if one wishes to introduce additional variables in quantum mechanics, to build the dBB (de Broglie-Bohm) theory [2, 3, 4]. Within Einsteinian relativity, we will see that the situation is more complex, and that it is not possible to define a probability current that takes consistent values in different Lorentz frames.
The essence of our discussion is very similar to that already given by Hardy [1], but here we emphasize the role of the probability current, instead of the “elements of reality”. In § 1, we first analyze the experiment within Galilean relativity, with absolute time. We show that, even in this simple case, the probability current exhibits unexpected contextual properties; in particular, it can change abruptly when the experimental conditions (the context) are slightly modified. In § 2 we discuss the same experiment within Einsteinian relativity, and show that the properties of the probability current are even more surprising: when the same experiment is described in a different Lorentz reference frame, the probability current may look very different; it may even exhibit (almost) discontinuous variations that have very little to do with those of the current of a classical fluid. This puts severe constraints on the construction of theories with additional variables.
Somewhat related considerations have been developed by Gisin [5], who emphasizes that even nonlocal deterministic theories cannot reproduce the results of quantum mechanics. Berndl et al. [6] propose a discussion that is also similar to ours; they put the emphasis on the absence of relativistic quantum equilibrium and on the possibilities offered by a multitime dBB formalism, leading to the introduction of synchronized $N$-path for $N$ noninteracting particles. Here we remain within the scope of standard quantum mechanics to discuss, for spinless particles, peculiar properties of the probability current.
1 Galilean relativity
Following Hardy [1], we consider two interferometers $\mathcal{A}$ and $\mathcal{B}$ (figure 1). The particles crossing these devices are entangled since, if the particles fly across the arms $a_{2}$ and $b_{2}$ in the middle of the figure, they interact and annihilate each other by some process.
Figure 1: Each of the two interferometers $\mathcal{A}$ and $\mathcal{B}$ receives a particle. The oval symbolizes an interaction process that destroys the two particles if they both take the middle paths. This process creates an entanglement between the two particles.
We denote by $|a_{0}\rangle$ and $|b_{0}\rangle$ the states of the particles entering the interferometers, by $|a_{1}\rangle$ and $|a_{2}\rangle$ (first particle), $|b_{1}\rangle$ and $|b_{2}\rangle$ (second particle), their quantum states when they propagate inside the inteferometers, as indicated in figure 1. Finally $|A_{1}\rangle$ and $|A_{2}\rangle$, $|B_{1}\rangle$ and $|B_{2}\rangle$ are the quantum states of each particle after it crosses the output beam splitter. The effect of these output beam splitters is described by the substitutions [7]:
[…]
and similar substitutions for the interferometer $\mathcal{B}$, as well as for the input beam splitters.
1.1 State vectors
Just after crossing the two first beam splitters, the system is in state:
$|a_{0}\rangle|b_{0}\rangle\Rightarrow\frac{1}{2}\left[i|a_{1}\rangle+|a_{2}\rangle\right]\left[i|b_{1}\rangle+|b_{2}\rangle\right]$ (2)
Suppressing the $|a_{2}\rangle|b_{2}\rangle$ component leads to the entangled state of the two particles:
$|\psi\rangle=\frac{1}{\sqrt{3}}\left[-|a_{1}\rangle|b_{1}\rangle+i|a_{1}\rangle|b_{2}\rangle+i|a_{2}\rangle|b_{1}\rangle\right]$ (3)
If one extends the arms of interferometer $\mathcal{B}$ a shown in figure 2, and if one is interested in the state of the system before the second particle reaches its output beam splitter (or if this beam splitter is removed), one can use relation (1) to obtain the state :
$|\psi_{A}^{\prime}\rangle=\frac{1}{\sqrt{6}}\left[-2\,|A_{1}\rangle|b_{1}\rangle+i\,|A_{1}\rangle|b_{2}\rangle-\,|A_{2}\rangle|b_{2}\rangle\right]$ (4a)
We note the absence of a component in $|A_{2}\rangle|b_{1}\rangle$. If, instead, we consider the situation of figure 3 and take into account only the output beam splitter of interferometer $\mathcal{A}$, we obtain the state:
$|\psi_{B}^{\prime}\rangle=\frac{1}{\sqrt{6}}\left[-2\,|a_{1}\rangle|B_{1}\rangle+i\,|a_{2}\rangle|B_{1}\rangle-|a_{2}\rangle|B_{2}\rangle\right]$ (4b)
from which the component in $|a_{1}\rangle|B_{2}\rangle$ has now disappeared.
Finally, when both output beam splitters are taken into account, the system is described by the ket:
$|\psi^{\prime\prime}\rangle=\frac{1}{\sqrt{12}}\Big{[}-3\,|A_{1}\rangle|B_{1}\rangle-i\,|A_{1}\rangle|B_{2}\rangle-i|A_{2}\rangle|B_{1}\rangle-|A_{2}\rangle|B_{2}\rangle\Big{]}$ (5)
which maintains a non-zero component on all possible output channels. This is a consequence of the suppression of the component in $|a_{2}\rangle|b_{2}\rangle$ in (3): if it is retained, applying relation (2) twice to this state leads to the simple relation:
$|a_{0}\rangle|b_{0}\rangle\Rightarrow i|A_{1}\rangle|B_{1}\rangle$ (6)
Therefore, if the particles do not interact, they all exit through channel $|A_{1}\rangle|B_{1}\rangle$.
1.2 Probability currents and trajectories
We now consider situations where two particles, described by wave packets, enter simultaneously the two interferometers $\mathcal{A}$ and $\mathcal{B}$. The Schrödinger probability current $\bm{J}$ has 6 components in configuration space (3 containing the derivatives of the wave function with respect to the coordinates of the first particle, 3 those for the other particle), and provides the direction along which the pair of the two wave packets moves in this space. One can then build continuous trajectories in configuration space that are tangent to $\bm{J}$ at each of their points. Since they are nothing but pairs of trajectories in usual 3D space, and in order to avoid confusion with ordinary trajectories, we will call them “bi-trajectories”. In general (if the wave function is a not product of single particle wave functions), the current depends on the positions of both particles; the 3D trajectory of each particle depends on the position of the other on its own trajectory.
The properties of bi-trajectories may look significantly different from those of a 3D trajectory for a single particle. While in the latter case, a 3D no-crossing rule limits the possible patterns, for bi-trajectories the no-crossing rule applies only in the configuration space, so that crossing in ordinary space are indeed possible. For instance, several different trajectories my start from the same point in interferometer $\mathcal{A}$, provided they are associated with different trajectories in the other interferometer. Moreover, the direction of each 3D trajectory may be affected by the insertion of a beam splitter recombining two wave functions in the other interferometer. Of course these properties occur only in the presence of entanglement, which introduces quantum nonlocal effects.
With wave packets, the probabilities of two-particle measurement results can be obtained as the time integral of the 6D flux of the probability current through the input surface of a detector. This operation should nevertheless be performed only with the values of the current at a given time; one should not use the values of the unperturbed probability current at different times to obtain correlation functions for measurement at different times. To correctly obtain these functions, one must take into account the effect of the first measurement on the state vector, which changes the probability current in both interferometers (even if the measurement is performed in only one of them). In Bohmian theory for instance, the calculation of time correlation functions requires that the perturbation of the first measurement on the evolution of the Bohmian positions is included, before the probability of the second measurement is calculated [9].
Inside the interferometers, the possible trajectories associated with the setup shown in figure 1 are obtained from the state vector of equation (3), while the state written in (5) determines the trajectories after the particles have crossed the output beam splitters. No pair of trajectories follows the path $a_{2}b_{2}$, since the corresponding component of the state vector vanishes: if one 3D trajectory takes the inner path, the other has to take the outer path. Whatever the detailed configurations of the bi-trajectories is, relation (5) indicates that they reach the 4 output channels, since they indicate the existence of a probability current that brings the probability into these regions.
1.3 Different configurations
Figure 2: A different version of the experiment where the crossing of the output beam splitter occurs first in interferometer $\mathcal{A}$. The lines of the probability current in configuration space (bi-trajectories) are then modified in this interferometer. Moreover, if the bi-trajectory follows $b_{1}$ in the lower interferometer, it is certain that it goes to $A_{1}$ in the upper interferometer. In other words, the only bi-trajectories that may exit through $A_{2}$ must also pass through $b_{2}$
Now, consider the situation shown in figure 2, where the output of beam splitter of $\mathcal{B}$ has been pushed to the right. The particle in interferometer $\mathcal{A}$ then crosses its output beam splitter before the other; in the intermediate time, the state of the system is given by relation (4a). This expression shows that, if the trajectory is such that the particle in interferometer $\mathcal{B}$ follows path $b_{1}$, the particle in interferometer $\mathcal{A}$ necessarily exits through $A_{1}$. The only the bi-trajectories where one particle reaches $A_{2}$ must have the other following path $b_{2}$.
Figure 3: A third version of the experiment where the crossing of the output beam splitter occurs first in interferometer $\mathcal{B}$. After this crossing, if the bi-trajectory follows path $a_{1}$ in interferometer $\mathcal{A}$, it is certain that the other component reaches $B_{1}$. The only bi-trajectories that may reach $B_{2}$ must also pass through $a_{2}$.
In figure 3, we consider the symmetric situation where it is the beam splitter of interferometer $\mathcal{A}$ that has been pushed to the right. The state of the system is given by relation (4b) and, in this case, the only bi-trajectories that may reach exit $B_{2}$ have to pass through $a_{2}$.
1.4 Discussion
According to relation (3) the two particles cannot be found simultaneously in $a_{2}$ and $b_{2}$, which shows that no bi-trajectory takes this path. Relation (5) indicates that some bi-trajectories reach $A_{2}$ and $B_{2}$ (more precisely, 1/12 of the trajectories). But, in § 1.3, we concluded in one case that the only trajectories reaching $A_{2}$ have to pass trough $b_{2}$, and in the other case that the only trajectories reaching $B_{2}$ have to pass through $a_{2}$; combined together, these two statements seem to indicate that the trajectories reaching $A_{2}$ and $B_{2}$ have to pass through $a_{2}$ and $b_{2}$, which is impossible as we just saw.
This apparent contradiction can be lifted by remarking that we have combined statements that are valid under incompatible experimental configurations, those illustrated in figures 2 and 3. As a consequence, there is no reason why the probability current should have the same configuration in both cases. It is actually context dependent, and what matters is the position of beam splitters and mirrors in the whole apparatus.
The mechanism behind this sensitivity of the probability current to the whole experimental setup is nonlocal in ordinary 3D space. When the particle in $\mathcal{B}$ crosses its beam splitter, components of the wave function that were spatially disjoint (in terms of the positions of that particle) are made to overlap, which introduces interference terms for the wave function guiding the particle in $\mathcal{A}$. This nonlocal effect is directly seen on the expressions (4a) to (5) of the quantum state of the physical system.
An striking illustration of this nonlocality of the probability current is the appearance of discontinuities when the experimental situations are modified, more precisely very rapid changes of the current lines when one element of the experimental device is slightly changed. When the position of one beam splitter is moved in order to progressively change the setup from that of figure 2 to that of figure 3, the probability current undergoes a sudden variation when the situation of figure 2 is crossed. Initially, all bi-trajectories reaching $A_{2}$ and $B_{2}$ pass through $a_{1}$ and $b_{2}$, while at the crossing they jump discontinuously to $a_{2}$ and $b_{1}$ (in dBB theory, the evolution of the Bohmian position necessarily exhibits the same discontinuity).
More precisely, when wave packets of finite length $\ell$ are crossing the output beam splitters, the state vector may actually be a superposition of states of the form (4) and (5). The rapid change of the pattern of the current lines should then take place when the length $L$ of one interferometer varies by $\ell$ around the length of the other interferometer. So, in fact no real mathematical discontinuity should occur, but just a very rapid change between two regimes of the current lines for a small variation of $\ell$. Since $L$ may be arbitrarily large, the crossover may seen as a quasi-discontinuity at the scale of $L$.
Hardy has also discussed this situation in terms of “empty waves” and their nonlocal effects. In any theory where localized particles are guided by a wave, the very possibility of observing pairs of particles exiting through outputs $A_{2}$ and $B_{2}$ is indeed paradoxical: on one hand, such an event is possible only if the particles interact by taking the paths $a_{2}$ and $b_{2}$, as seen by comparing relations (5) and (6); but, on the other hand, they are supposed to disappear if they follow these trajectories !
Ref. [10] discusses another version of the experiment, where the oval in figure 1 symbolizes a dephasing interaction, instead of the mutual annihilation of the two particles. This variant of the experiment may be relevant in the context of the detection of quantum effects induced by gravity [11]. Interestingly, the two versions of the experiment lead to very similar conclusions.
2 Einsteinian relativity
We now discuss how, in special relativity, the sensitivity to the context induces a sensitivity to the Lorentzian reference frame, preventing a Lorentz invariant definition of the probability current. We first specify the properties that we assume to be valid within the quantum theory under discussion.
2.1 General properties
Within a relativistic quantum theory, the measurement results (indications of the pointers of the apparatuses) and their probabilities are relativistic scalars. If for instance a Stern-Gerlach experiment is observed from different Lorentz frames, even if the various observers see a different motion of the atoms and of the magnet, they all agree that the possible measurement results are ±/2plus-or-minusPlanck-constant-over-$\pm\hbar/2$, as well as on the values of their probabilities. This also applies to correlation measurements, provided of course the space-time coordinates of the events are properly modified.
Within a relativistic quantum theory, we consider a physical system made of a fixed number $N$ of massive particles, and we assume that the theory contains a local conservation law of the probability. This law involves a probability current defined in configuration space, taking different value in different Lorentz frames; nevertheless, in each frame, a local law of conservation of the probability can be written with its own current. This current is closely related to the probabilities of measurement results since, with wave packets, the flux of this current (integrated over time) through appropriate surfaces (for instance the input surfaces of detectors) allows one to calculate probabilities, i.e. relativistic invariants.
For a system of two particles, the flux has to be calculated at a given time through a surface in a 6D space, in other words through a pair or surfaces in ordinary 3D space. As above, the probability current $\bm{J}$ allows one to define bi-trajectories, which are tangent to $\bm{J}$ at each point of configuration space. One can then obtain the probabilities of measurement results by making statistics over the trajectories, which are nothing but statistics over their starting point (with a weight given by the initial quantum probability of presence). This provides a view of the theory where the random character does not appear during measurements, but only in the value of the initial positions of the trajectories.
The current can be used in each Lorentzian frame to obtain the probability of any measurement performed at a given time in this frame. We will therefore use it for simultaneous measurements performed with two detectors sitting at any place in each interferometer (before or after the beam splitters), but not to derive correlations between results of measurements performed at different times – since this would require to take into account the perturbation of the first measurement on the current [9]. This necessity can be seen as a consequence of the von Neumann projection postulate, or of the entanglement of the system with the measurement apparatus (for instance, within dBB theory, the Bohmian positions within this apparatus play an important role [12]).
Simultaneous measurements in one Lorentz frame appear as time shifted measurements in another frame. We will deduce the properties of the probability current in each frame from probabilities of simultaneous measurements in the same frame, and then use the fact that probabilities are relativistic scalar to obtain them from calculations in a different frame.
2.2 Different views in different Lorentzian frames
We now consider an double interference experiment which, in the Lorentz frame of the laboratory $\mathcal{L}_{0}$, corresponds to the scheme shown in figure 1. In another Lorentz frame, the same experiment is seen differently. For instance, figure 4 shows a time-space diagram describing the experiment in a frame $\mathcal{L}_{1}$ moving with a positive velocity $v$ along axis $Ox$. The output beam splitter is then crossed in interferometer $\mathcal{A}$ before than it is in interferometer $\mathcal{B}$. Thus, in this Lorentz frame, one can consider a simultaneous measurement of the output channel in interferometer $\mathcal{A}$ while the other particle is still in interferometer $\mathcal{B}$; the probability of any result can then be obtained by time integration of the current in this frame. To obtain this probability, we consider the same measurement performed in the laboratory frame, which becomes an time-shifted measurement experiment and corresponds to the configuration of figure 2. The probability of this two time measurement in the laboratory frame is then obtained by using the standard Schrödinger theory and the expression (4a) of the state vector. Finally, we use the fact that probabilities are relativistic scalars to obtain the probability of the measurement as seen in the moving frame $\mathcal{L}_{1}$, and from this infer properties of the probability current and trajectories in this frame.
For the same reason as in § 1, if the detectors are inserted in channel $A_{2}$ and $b_{1}$, the probability vanishes, but does not if they are inserted in channel $A_{2}$ and $b_{2}$. All bi-trajectories that reach $A_{2}$ therefore have to pass through $b_{2}$. If we extend the (unperturbed) trajectories by continuity we see that, in Lorentz frame $\mathcal{L}_{1}$, all bi-trajectories reaching outputs $A_{2}$ and $B_{2}$ necessarily pass through $b_{2}$. Since, as above, expression (3) shows that no trajectory passes through channels $a_{2}$ and $b_{2}$, we conclude that these trajectories necessarily pass through $a_{1}$ and $b_{2}$.
[…]
Figure 4: This figure shows the experiment of figure 1 in another Lorentz frame, moving with a positive velocity $v$ in the direction of the vertical $Ox$ axis in the figure. The particles are assumed to have a non-zero mass. In interferometer $\mathcal{A}$, in some sections the propagation is parallel to $Ox$ to axis, which increases the velocity; in interferometer $\mathcal{B}$ it is antiparallel, which decreases the velocity. As a consequence, the crossing of the output beam splitter occurs first in interferometer $\mathcal{A}$ as in figure 2.
3 Conclusion
We therefore have a situation where, while the various Lorentz observers agree on the values of the probabilities of all measurement results, they may disagree on the path taken by the probability to reach the detectors. Curiously, while they agree when they calculate the flux of the probability current (more precisely, its integral over time), they disagree on the lines of current that lead to this flux. The major difference of points of view occurs between observers with positive velocity $v$ and those with negative velocity. When $v$ vanishes, a discontinuous change occurs, which is reminiscent of the discontinuity taking place in Galilean relativity when the experimental setup is changed. Nevertheless, a discontinuity of the description of the same experiment by different observers raises a more serious problem, since it cannot be solved by invoking the Bohrian wholeness of the experimental apparatus. It therefore seems difficult to attribute physical reality to the probability current, and consequently to the probability fluid itself.
All our reasoning took place within standard quantum mechanics, but of course it has consequences if one uses the dBB theory. A pointed out in Ref. [6], the so called quantum equilibrium condition of this theory cannot be relativistically invariant. One possibility is to decide that the theory requires the choice of a privileged frame, for instance that of the laboratory (but, even with this rule, problems arise when one considers experimental setups containing rapidly moving components [13, 14]) or the frame of the cosmic microwave radiation.
But one can also take another point of view. The dBB theory achieves two main results: first, it unifies the dynamics, considering that measurements are just ordinary quantum interaction processes, and eliminating the need for any specific postulate (von Neumann projection, Heisenberg cut, etc.); second, it provides a description of reality (an ontology) expressed in terms of point positions of particles. Clearly, the second result is incompatible with a completely relativistic theory [^1] In the context of field theory, Bell [15] discusses why local occupation numbers (beables) for a fermionic field cannot provide a fully relativistic view of reality either., where no Lorentz frame is privileged. But, even if one gives up the description of reality with point particles, one can still preserve the first result: the positions are then considered just as an indicator pointing to the branch of the state vector that is active, and eliminates all the empty waves. In this way, the unification of the dynamics is preserved, and the closest description of reality is obtained in terms of the active branch of the state vector. True, the indicator that determines this active branch is not relativistically invariant but, since it appears as a purely mathematical variable in the dynamics, this does not necessarily ruin the approach.
In conclusion, and whether or not a unified dynamics is desirable in quantum mechanics, it is clear that even blurred relativistic trajectories seem to be incompatible with this theory – this remains true if they are blurred more than what is requested by the Heisenberg uncertainty relations. Bohr went even further [16] by claiming that, due to the non-zero value of the Planck constant, “the space-time description and the claims of causality are of complementary nature”, and therefore incompatible; this should indeed forbid a description of the motion of the probability fluid in space-time. Discontinuities in the trajectories taking place when the Lorentz reference frame is changed may be seen as a particular case of application of this general statement.